# Harmonia: Algorithm-Hardware Co-Design for Memory- and Compute-Efficient BFP-based LLM Inference

### v1

@@ Abstract
Large Language Models (LLMs) are powerful but costly to run in terms of memory and compute. Quantization is particularly effective since it reduces both model size and compute cost. A widely adopted configuration uses integer (INT) weights with floating-point (FP) activations to preserve accuracy. To further reduce the overhead of FP arithmetic, prior works convert FP activations in linear layers into block floating point (BFP), where a group of activations shares a single exponent. However, prior works failed to extend BFP to attention layers due to significant accuracy loss, leading to suboptimal efficiency.
To address this issue, we propose Harmonia, an algorithm-hardware co-design framework that supports an all-layer BFP format and integrates a configurable hardware design. First, to enable all-layer BFP-activations, we systematically explore BFP configurations to achieve a better trade-off between model accuracy and activation compression. Second, to further reduce KV-cache related storage and computation in attention layers, we propose an asymmetric bit-allocation strategy, combined with a hybrid offline-online outlier smoothing technique. Based on this, we aggressively compress the KV-cache from FP16 to 4b-mantissa BFP, with an average loss of only 0.3%. Third, to fully exploit the benefits of all-layer BFP-activations, we propose a suite of dedicated hardware designs for Harmonia. Specifically, we propose a reconfigurable PE unit supporting mixed data formats (i.e., BFP-INT, BFP-BFP) and mixed data precisions, a real-time FP16-to-BFP converter for on-the-fly activation compression, and a flexible tiling-aware dataflow to reduce external memory traffic. We evaluate Harmonia on general matrix multiplication (GEMM) in both linear layers and attention layers across eight widely used LLMs. Compared with prior works, experimental results show that Harmonia achieves 3.84× (up to 5.05×) higher area efficiency, 2.03× (up to 3.90×) better energy efficiency, and 3.08× (up to 4.62×) speedup on average, respectively.
@@ I Introduction
Large Language Models (LLMs) [59], [60], [26], [68], [1], [51], [10] have emerged as a transformative foundation in artificial intelligence, demonstrating remarkable proficiency across a wide range of natural language processing tasks, including recommendation systems, chatbots, and personal agents. Guided by the scaling law [32], the most advanced LLMs’ parameters have grown to hundreds of billions, posing substantial demands on both memory and computational resources. For instance, deploying a Llama-3.1-405B model, which occupies around 800GB of memory, requires at least 10 high-end H100 GPUs. Additionally, during the generation phase, the auto-regressive nature of LLMs necessitates recurrent retrieval of a large number of previously generated tokens, resulting in a memory bandwidth bottleneck [37], [2], [34], [56], which severely impacts the performance and efficiency of model inference. These challenges emphasize an urgent need to improve the efficiency of LLM deployment.
@@
Quantization is an effective approach for efficient model deployment. [72], [63], [3], [65], [41], [44], [58] quantize both floating-point (FP) weights and FP activations to low-precision integers. While weights can be statically quantized offline after training, the dynamic nature of activations necessitates runtime quantization, introducing additional latency and hardware costs [24], [36], [69], [33]. As a practical alternative, the weight-only quantization scheme [17], [13], [62], [40], [47], [11], [35], [12], [53] where only weights are quantized to low bits (e.g., INT4), while activations remain in FP format (e.g., FP16), has been widely adopted. It can effectively reduce the model size while preserving the model’s linguistic capabilities. However, this scheme still involves costly FP-related operations (e.g., exponent alignment and normalization) and FP-format storage [25], leading to significant hardware overhead.
@@
To alleviate this issue, FIGNA [25] and Anda [16] propose to convert FP activations into the block floating point (BFP) format, where all elements within a block share a common exponent. This shared-exponent feature eliminates the need for costly per-element exponent alignment and normalization, enabling more efficient BFP-related multiplication and accumulation (MAC) on hardware.
@@
However, these designs have two major limitations. First, they fail to extend the BFP format to FP activations in attention layers, which constitute a critical component of LLM architectures. As shown in Fig. 1(a), the dominant computation shifts between linear and attention layers depending on the input sequence length and model configuration. In short-context scenarios [6], [38], linear layers prevail across models of different scales, whereas in long-context applications [4], [70], [29], attention layers rapidly become the primary computational bottleneck as sequence length increases. Consequently, optimizing both layers is equally critical to achieving efficient and flexible adaptation across diverse workloads. Since both designs leave attention activations to be processed by conventional FP-based units, the computation cost still leaves substantial room for improvement.
@@
Second, both designs retain the Key-Value (KV) cache in FP16 format, overlooking the frequent external memory accesses (EMA) it incurs. The KV cache stores intermediate attention keys and values, and can occupy over 90% of the total memory footprint when the sequence is sufficiently long [34]. As shown in Fig. 1(b), its size scales linearly with sequence length, resulting in substantial EMA overhead. This excessive traffic becomes a major bottleneck during the decoding stage [23], [37], [22], where computation is largely memory-bound and inference throughput is severely constrained. Therefore, maintaining the KV cache in full precision yields no tangible benefit in reducing EMA power or bandwidth pressure.
@@
To overcome these limitations, we propose Harmonia, an algorithm-hardware co-design framework that incorporates a unified BFP representation across activations in both linear and attention layers, supported by a configurable compute core capable of executing both BFP-INT and BFP-BFP operations. We begin by empirically exploring the feasibility of extending the BFP format to all-layer activations. Through experiments, we identify a BFP configuration that balances model accuracy and activation compression. Building upon this basis, we further propose an asymmetric bit-allocation strategy and a hybrid offline-online outlier smoothing technique to compress the KV cache from FP16 to 4-bit mantissas without noticeable accuracy loss (only 0.3% on average).
@@
  • We extend the BFP format to all activations in linear and attention layers, adopting a well-selected configuration with a group size of 32 and 8-bit mantissas for all activations except the KV cache.
@@
Figure 2: Illustration of auto-regressive LLM architecture.
@@ II Background and Motivation
This section first outlines the LLM inference process, then introduces the existing quantization techniques. After that, we further analyze the limitations of existing methods, which reveal opportunities and motivate our design.
II-A LLM Inference Process
@@
In terms of computation, both linear and attention layers in LLMs are fundamentally composed of GEMM in the prefill stage and GEMV in the decode stage. Supposing the hidden dimension of the model is $d$, the lengths of the input prompt and the generated tokens are $n$ and $l$ separately, the operations in linear layers scale with $(n+l)d^{2}$, whereas those in attention layers scale with $d(n^{2}+\frac{1}{2}l^{2}+(n-1)l)$. Importantly, since the sequence length in real-world LLM applications varies from thousands to millions [14], [5], while $d$ remains constant after training, the computational bottleneck shifts between linear and attention layers depending on the workload. In terms of memory, the footprint is dominated by two major components: weights and the KV cache, either of which can become the limiting factor under different scenarios [56], [37].
Therefore, alleviating the computational overhead in both linear and attention layers, as well as reducing accesses to weights and the KV cache stored in external memory, is critical to improving inference efficiency.
II-B Quantization Techniques
Quantization is an effective technique that converts high-precision values to low-bit formats, reducing computation and memory costs. In particular, quantizing weights to integer formats offline has been shown to be highly effective. However, activation quantization is more challenging due to its wider value distributions and the presence of outliers, which can severely degrade accuracy. Moreover, it requires dynamically computing scaling factors and offsets to adapt to the runtime value variations. This introduces additional hardware into existing pipelines [24] and sometimes requires system-level support [43], [72] to fully unlock its benefits.
While the INT-weight and FP-activation scheme can greatly reduce the memory cost for LLM inference, it cannot avoid expensive FP-based MAC operations, which require exponent alignment and normalization. To mitigate this issue, recent works convert FP activations in linear layers to the BFP format, which shares the same exponent within a group. It achieves a balance between the dynamic range of FP formats and the simplicity of INT-based MAC operations.
@@ II-B Quantization Techniques
Fig. 3 shows the steps to convert an FP16 vector to BFP: (1) partition the vector into groups, (2) extract the largest exponent in each group as the shared exponent, and (3) right-shift and truncate mantissas based on exponent differences. Evidently, the group size and mantissa precision directly determine the numerical error introduced by this process. In general, smaller groups with sufficient mantissa bits yield lower error, while larger groups and aggressive truncation increase the error, which can degrade model accuracy. Thus, selecting an appropriate BFP configuration to balance model accuracy and efficiency remains a non-trivial challenge.
Figure 3: Illustration of the FP16-to-BFP conversion with a group size of 3 and an 8-bit mantissa.
@@
To mitigate the impact of numerical errors on model accuracy, one feasible strategy is BFP-aware training [67],[9], [50], [18], which directly adopts the BFP format during training. However, LLM training is prohibitively expensive [8], [21], [28], [60], requiring massive computation and time, which limits its practicality. Instead, another line of work introduces BFP conversion at inference time. FIGNA [25] adopts a conservative strategy by analytically relating mantissa bit width to numerical precision and extending the mantissa to match FP-based results. Although this preserves accuracy, storing activations in FP16 format eliminates both memory and bandwidth savings. Anda [16] employs a customized search algorithm to allocate variable mantissa bit widths to different types of activations in linear layers under bounded accuracy loss, aiming to minimize total bit operations.
@@ II-C Opportunities and Challenges
The results indicate that model accuracy is highly sensitive to the mantissa bit width. Each subplot of Fig. 4 demonstrates that different models exhibit varying levels of accuracy degradation under the same BFP configuration, but they share a consistent trend: accuracy decreases as the mantissa bit width shrinks. Notably, when the bit width drops below 8 bits, accuracy declines sharply. Using 8-bit mantissas keeps the accuracy degradation around 1.5% for most models, which we consider an acceptable loss. Therefore, all subsequent analyses and experiments adopt 8-bit mantissas for activations except those in the KV cache.
@@ III Harmonia Algorithm
This section presents the algorithmic foundation of Harmonia. Based on the experimental results in Sec. II-C, we set the group size to 32, adopt a 5-bit shared exponent per group, and use an 8-bit mantissa for activations except the KV cache. In this section, we begin by introducing a dynamic BFP conversion framework, which converts activations to the BFP format during inference. To mitigate the accuracy loss caused by compressing the KV cache to 4-bit mantissas, we propose two key optimizations: (1) an initial-local asymmetric bit allocation strategy and (2) an offline–online hybrid outlier smoothing technique. Finally, we analyze the hardware benefits of BFP-based conversion, which motivate the unified and efficient architecture of Harmonia.
@@ III-C Offline-Online Hybrid Outlier Smoothing
III-D Unified Integer-based BFP-INT and BFP-BFP MACs
Building on the above algorithmic optimizations, we enable the conversion of activations in both linear and attention layers to the BFP format during inference. This transformation lays the groundwork for a unified hardware execution of BFP-INT and BFP-BFP MACs on a single compute core. In conventional FP-based inner-product computation, partial products carry different exponents, requiring dynamic alignment through a costly arbitrary-bit barrel shifter in each PE. In contrast, the BFP format enforces a shared exponent within each group, allowing direct integer accumulation without per-element alignment. The accumulated result is then normalized once per group, amortizing the normalization overhead. This numerical format conversion simplifies the datapath and enhances computational efficiency.
To intuitively demonstrate the hardware benefits introduced by Harmonia’s algorithms, Fig. 11 illustrates the computation flows of the linear and attention layers across four representative designs. Fig. 11(a) shows that the dedicated FP-INT engine removes the FP dequantization step required in GPUs but still suffers from costly alignment and normalization inherent in FP-based calculation. As shown in Fig. 11(b) and (c), both FIGNA [25] and Anda [16] employ FP-to-BFP conversion to perform integer-based MACs. FIGNA retains FP16 activations and converts them to extended-mantissa BFPs before computation, incurring extra access and conversion overhead without reducing memory cost. Anda stores activations in variable mantissa bit widths to cut memory traffic, but its bit-serial design adds timing and hardware complexity. However, all three designs still rely on an additional FP-FP engine shown in Fig. 11(d) for attention layers, diminishing overall performance gains when attention layers dominate computational workload in long-context scenarios.
Fig. 11(e) illustrates the key advantages of Harmonia over prior designs. First, both linear and attention operations can be executed on a single reconfigurable core (detailed in Sec. IV), enabling efficient resource sharing and high hardware utilization. Second, truncated-mantissa BFP representation of activations lowers storage overhead. Finally, to mitigate the KV cache growth during generation, Harmonia compresses it to 31.25% of its original size, substantially reducing storage footprint and memory access overhead.
@@ IV-C Real-time BFP Converter
Specifically, for V-type activations grouped per channel with a group size of 32, as shown in the upper left of Fig. 14(c), all eight results within a column (e.g., $C_{0,0}$-$C_{7,0}$) are fed into a comparator tree simultaneously to identify the local maximum exponent. Since each group contains 32 elements, four output loops are required to obtain the final maximum exponent. The 32 elements of each group are then divided into four 8-element subgroups, which are forwarded to the aligner along with the derived maximum exponent. Each step processes 8 mantissas in parallel to produce the converted results.
@@
Figure 15: Configurable output dataflows: column-major and row-major modes.
@@ IV-D Tiling-aware Dataflow Design
TABLE I: Perplexity on WikiText2. Lower is better. BFP(x) denotes an x-bit mantissa precision.

### v2

@@ Abstract
Large Language Models (LLMs) are powerful but incur high memory and computation costs. Weight-only quantization is an effective solution with integer (INT) weights and floating-point (FP) activations being adopted to preserve accuracy. To further reduce the overhead of FP-INT arithmetic, prior works convert FP activations in linear layers into block floating point (BFP), where a group of activations shares a single exponent. However, prior works failed to extend BFP to attention layers due to significant accuracy loss, leading to suboptimal efficiency. To address this issue, we propose Harmonia, an algorithm-hardware co-design framework that enables all-layer BFP activations with a configurable hardware architecture. First, we systematically explore BFP configurations to achieve a better trade-off between accuracy and compression rate across all layers. Second, to further reduce KV-cache related storage and computation in attention layers, we propose an asymmetric bit-allocation strategy, combined with a hybrid offline-online outlier smoothing technique. Based on this, we aggressively compress the KV-cache from FP16 to 4b-mantissa BFP, with an average loss of only 0.3% on LongBench benchmark. Third, to fully exploit the benefits of all-layer BFP-activations, we propose a suite of dedicated hardware designs for Harmonia. Specifically, we propose a reconfigurable PE unit supporting mixed data formats (i.e., BFP-INT, BFP-BFP) and mixed data precisions, a real-time FP16-to-BFP converter for on-the-fly activation compression, and a flexible tiling-aware dataflow to reduce external memory traffic. We evaluate Harmonia in both linear layers and attention layers across eight widely used LLMs. Compared with prior works, experimental results show that Harmonia achieves 3.84× (up to 5.05×) higher area efficiency, 2.03× (up to 3.90×) better energy efficiency, and 3.08× (up to 4.62×) speedup on average, respectively.
@@ I Introduction
Large Language Models (LLMs) [51], [52], [23], [59], [1], [2], [38] have emerged as a transformative foundation in artificial intelligence, demonstrating remarkable proficiency across a wide range of natural language processing tasks, including recommendation systems, chatbots, and personal agents. Guided by the scaling law [28], the most advanced LLMs’ parameters have grown to hundreds of billions, posing substantial demands on both memory and computational resources. For instance, deploying a Llama-3.1-405B model, which occupies around 800GB of memory, requires at least 10 high-end H100 GPUs. Additionally, during the generation phase, the auto-regressive nature of LLMs necessitates recurrent retrieval of a large number of previously generated tokens, resulting in a memory bandwidth bottleneck [33], [30], [48], which severely impacts the performance and efficiency of model inference. These challenges emphasize an urgent need to improve the efficiency of LLM deployment.
@@
Quantization is an effective approach for efficient model deployment. [63], [55], [3], [36], [39], [50] quantize both floating-point (FP) weights and FP activations to low-precision integers. While weights can be statically quantized offline after training, the dynamic nature of activations necessitates runtime quantization, introducing additional latency and hardware costs [21], [32], [60], [29]. As a practical alternative, the weight-only quantization scheme [14], [10], [54], [35], [31], [9] where only weights are quantized to low bits (e.g., INT4), while activations remain in FP format (e.g., FP16), has been widely adopted. It can effectively reduce the model size while preserving the model’s linguistic capabilities. To fully harness the computational gains from weight-quantized models, specialized execution units capable of processing FP-INT operands are typically required, as illustrated in Fig. 1(a). However, this scheme still involves costly FP-related operations (e.g., exponent alignment and normalization) and storage of FP-format activations, leading to significant hardware overhead.
@@
To alleviate this issue, as shown in Fig. 1(b) and (c), FIGNA [22] and Anda [13] propose to convert FP activations involved in linear layers into the block floating point (BFP) format, where all elements within a block share a common exponent. This shared-exponent feature eliminates the need for costly per-element exponent alignment and normalization, enabling more efficient BFP-related multiplication and accumulation (MAC) on hardware.
@@
However, these designs have two major limitations. First, they fail to extend the BFP format to FP activations in attention layers, which constitute a critical component of LLM architectures. As shown in Fig. 2(a), the dominant computation shifts between linear and attention layers depending on the input sequence length and model configuration. In short-context scenarios [5], [34], linear layers prevail across models of different scales, whereas in long-context applications [4], [61] attention layers rapidly become the primary computational bottleneck as sequence length increases. Consequently, optimizing both layers is equally critical to achieving efficient and flexible adaptation across diverse workloads. Since both designs leave attention activations to be processed by conventional FP-based units shown in Fig. 1(d), the specialized arithmetic units optimized for linear layers remain underutilized during attention execution, leading to poor area efficiency and fragmented hardware resources.
@@
Second, both designs retain the Key-Value (KV) cache in FP16 format, overlooking the frequent external memory accesses (EMA) it incurs. The KV cache stores intermediate attention keys and values, and can occupy over 90% of the total memory footprint when the sequence is sufficiently long [30]. As shown in Fig. 2(b), its size scales linearly with sequence length, resulting in substantial EMA overhead in long-context scenarios. This excessive traffic becomes a major bottleneck during the decoding stage [20], [33], [19], where computation is largely memory-bound and inference throughput is severely constrained. Therefore, maintaining the KV cache in full precision yields no tangible benefit in reducing EMA power or bandwidth pressure.
@@
To overcome these limitations, as illustrated in Fig. 1(e), we propose Harmonia, an algorithm-hardware co-design framework that incorporates a unified BFP representation across activations in both linear and attention layers, supported by a configurable compute core capable of executing both BFP-INT and BFP-BFP operations. We begin by empirically exploring the feasibility of extending the BFP format to all-layer activations. Through experiments, we identify a BFP configuration that balances model accuracy and activation compression. Building upon this basis, we further propose an asymmetric bit-allocation strategy and a hybrid offline-online outlier smoothing technique to compress the KV cache from FP16 to 4-bit mantissas without noticeable accuracy loss (only 0.3% on average).
@@
  • We employ a unified BFP format to represent activations across both linear and attention layers, adopting a well-selected configuration with a group size of 32 and 8-bit mantissas for all activations except the KV cache.
@@
Figure 3: Illustration of the weight-only quantized LLM architecture.
@@ II Background and Motivation
This section first outlines the inference paradigm of weight-only quantized LLMs, followed by an introduction to the BFP format and its existing applications. After that, we further analyze the limitations of prior works, which reveal opportunities and motivate our design.
II-A Weight-only Quantized LLMs
@@
To reduce the costly memory footprint of LLM deployment, quantization has been widely applied to weights in linear layers. In this scheme, weights are quantized to low bit widths while activations remain in FP format. Compared to quantizing both weights and activations, weight-only quantization can compress model parameters to ultra-low bit widths with negligible accuracy loss, facilitating efficient deployment of LLMs on resource-constrained devices in edge computing scenarios.
Despite its advantages, weight-only quantization exhibits notable limitations in both computation and memory aspects. In terms of computation, quantized linear layers introduce mixed-format FP-INT operations, while attention layers continue to rely on FP-FP operations. Both operation types involve expensive exponent alignment and normalization [22], which complicate hardware implementations and limit efficiency gains. In terms of memory, the KV cache is still maintained in FP format. As illustrated in Fig. 2, when the sequence length increases, the KV cache gradually dominates memory storage, overtaking model weights as the primary memory bottleneck during inference.
Consequently, optimizing FP activations across all layers emerges as a key opportunity for improving the overall efficiency of weight-only quantized LLMs.
II-B Block Floating Point
To reduce the computational and storage overhead of FP activations without significantly degrading model performance, Block Floating Point (BFP) [11] has been proposed as a balanced trade-off between the wide dynamic range of traditional FP formats and the simplicity of INT formats. Fig. 4(a) shows the steps to convert an FP16 vector to BFP: (1) partition the vector into groups, (2) extract the largest exponent in each group as the shared exponent, and (3) right-shift and truncate mantissas based on exponent differences. For comparison, Fig. 4(b) and Fig. 4(c) illustrate the numerical conversion procedures and correspongding area and power costs of block-wise (BW) INT quantization and microscaling (MX)[45] INT quantization, respectively. Owing to its simplified conversion process, BFP significantly reduces the overhead of online numerical conversion. Compared to BWINT and MXINT quantizers, the BFP converter reduces the area consumption by 8.23× and 1.52×, and achieves power savings of 8.55× and 1.60×, respectively.
@@ II-B Block Floating Point
The BFP format further provides flexibility by allowing different group size and mantissa bit widths, enabling a tunable trade-off between model accuracy and hardware efficiency. In general, smaller groups with sufficient mantissa bits yield lower error, while larger groups and more aggressive mantissa truncation increase the error, which can degrade model accuracy. Consequently, selecting an appropriate BFP configuration to balance model accuracy and efficiency remains a non-trivial challenge.
@@
To mitigate the impact of numerical errors on model accuracy, one feasible strategy is BFP-aware training [58],[8], [15], which directly adopts the BFP format during training. However, LLM training is prohibitively expensive [7], [18], [25], requiring massive computation and time, which limits its practicality. Instead, another line of work introduces BFP conversion at inference time. FIGNA [22] adopts a conservative strategy by analytically relating mantissa bit width to numerical precision and extending the mantissa to match FP-based results. Although this preserves accuracy, storing activations in FP16 format eliminates both memory and bandwidth savings. Anda [13] employs a customized search algorithm to allocate variable mantissa bit widths to different types of activations in linear layers under bounded accuracy loss, aiming to minimize total bit operations.
@@
Figure 4: Illustration of (a) the FP16-to-BFP numerical conversion procedure, (b) a comparison of the BWINT, MXINT, and BFP conversion processes, and (c) the normalized hardware area and power breakdown of the corresponding implementations.
@@ II-C Opportunities and Challenges
The results indicate that model accuracy is highly sensitive to the mantissa bit width. Each subplot of Fig. 5 demonstrates that different models exhibit varying levels of accuracy degradation under the same BFP configuration, but they share a consistent trend: accuracy decreases as the mantissa bit width shrinks. Notably, when the bit width drops below 8 bits, accuracy declines sharply. Using 8-bit mantissas keeps the accuracy degradation around 1% for most models, which we consider an acceptable loss. Therefore, all subsequent analyses and experiments adopt 8-bit mantissas for activations except those in the KV cache.
@@ III Harmonia Algorithm
This section presents the algorithmic foundation of Harmonia. Based on the experimental results in Sec. II-C, we set the group size to 32, adopt a 5-bit shared exponent per group, and use an 8-bit mantissa for activations except the KV cache. In this section, we begin by introducing a dynamic BFP conversion framework, which converts activations to the BFP format during inference. To mitigate the accuracy loss caused by compressing the KV cache to 4-bit mantissas, we propose two key optimizations: (1) an initial-local asymmetric bit allocation strategy and (2) an offline–online hybrid outlier smoothing technique.
@@ IV-C Real-time BFP Converter
Specifically, for V-type activations grouped per channel with a group size of 32, as shown in the upper left of Fig. 14(c), all eight results within a column (e.g., $C_{0,0}$-$C_{7,0}$) are fed into a comparator tree simultaneously to identify the local maximum exponent. Since each group contains 32 elements, four output loops are required to obtain the final maximum exponent. The 32 elements of each group are then divided into four 8-element subgroups, which are forwarded to the aligner along with the derived maximum exponent. Each step processes 8 mantissas in parallel to produce the converted results
@@ IV-D Tiling-aware Dataflow Design
Due to the limited capacity of on-chip SRAM, tiling strategies are widely adopted to efficiently support large-scale matrix multiplication. As illustrated in Fig. 15, consider multiplying a matrix A of size $M\times K$ with a matrix B of size $K\times N$, resulting in an output matrix C of size $M\times N$. When adopting dataflow 1 (column-first output flow) to compute matrix C, the total number of EMA for matrices A and B is $\frac{N}{n}\times(M\times K)+K\times N$. In contrast, using dataflow 2 (row-first output flow) results in the total number is $\frac{M}{m}\times(K\times N)+(M\times K)$. Taking the linear layer as an example, for a given model, $K$ and $N$ are fixed, while $m$ and $n$ are determined by the on-chip resource allocation and thus remain constant. However, in practical LLM workloads, the number of input tokens can vary substantially, resulting in a wide dynamic range of $M$. Consequently, the choice between the two dataflows directly impacts the EMA cost under different workloads.
Therefore, we design a flexible data generation flow (FDGF) controller that can switch between column-first and row-first dataflows through lightweight configuration. This controller manages the read address order of two data dispatchers from on-chip SRAM and coordinates on/off-chip data exchanges. Specifically, under the column-first dataflow, weights remain in on-chip SRAM until all activations are processed, after which the next weight column is fetched from off-chip memory. Conversely, in the row-first dataflow, activations stay on-chip until computation with all weights is completed before being replaced with new activations. This adaptability enables the architecture to select the most memory-efficient dataflow for a given workload, reducing costly EMA and further improving overall system efficiency.
Figure 15: Configurable output dataflows: column-first and row-first modes.
V Evaluation
TABLE I: Area and power breakdown of the harmonia architecture.
   Component Area(mm2) Power(mW)
   PE Array 1.4212 114.7610
   Data Dispatcher 0.0729 3.8890
   On-Chip Memory 1.8567 180.8730
   K-offset Generator 0.0129 0.4267
   BFP Converter 0.0009 0.0263
   Vector Unit 0.0027 0.1180
   Data Collector 0.0123 0.3240
   FDGF Controller 0.0016 0.1820
   Total 3.3812 300.6000
TABLE II: Perplexity on WikiText2. Lower is better. BFP(x) denotes an x-bit mantissa precision.